OpenStudy (anonymous):

Suppose that the event of picking a face card from a standard deck of 52 cards is considered a success. If a card is drawn at random and replaced, and this process is repeated 11 times, what is the probability that the card selected was a face card 6 times?

OpenStudy (anonymous):

A. 0.019 B. 0.024 C. 0.004 D. 0.976 E. 0.995

OpenStudy (kropot72):

There are 12 face cards in a standard deck of 52 cards. Therefore on each draw the probability of drawing a face card is 12/52. The probability of drawing a face card 6 times in 11 draws is found as follows: \[P(6\ face\ cards)=\frac{11!}{6!5!}\times(\frac{12}{52})^{6} \times(\frac{40}{52})^{5}=you\ can\ calculate\]

OpenStudy (anonymous):

thank you so much!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

whats the answer